# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### v2

@@ Appendix C Experimental details
In our Atari experiments, we use the hyperparameters reported in Agarwal et al. [2020], and we do not tune the hyperparameter for our proposed method. The original dataset contains $5$ runs of DQN replay data. Each run of DQN replay data contains $50$ datasets, and each dataset contains $1$ million transitions. Due to the memory constraint, we can not load the entire data. As a result, for each training epoch, we select $5$ datasets randomly, subsample a total of $500$k transitions from the $5$ selected datasets, and perform $100$k updates using the $500$k transitions.
Figure 3 show the result with clipped losses and unclipped losses. Log-loss consistently outperforms other losses.

### v5

@@ Appendix C Experimental details
In our Atari experiments, we use the hyperparameters reported in Agarwal et al. (2020), and we do not tune the hyperparameter for our proposed method. The original dataset contains $5$ runs of DQN replay data. Each run of DQN replay data contains $50$ datasets, and each dataset contains $1$ million transitions. Due to the memory constraint, we can not load the entire data. As a result, for each training epoch, we select $5$ datasets randomly, subsample a total of $500$k transitions from the $5$ selected datasets, and perform $100$k updates using the $500$k transitions.
Figure 4 show the result with clipped losses and unclipped losses. Log-loss consistently outperforms the DQN variants other than C51.